OpenStudy (salina):

Ofelia is taking ten shots in the intramural free throw shooting competition.How many sequences of makes and misses are there that result in her making eight shots and missing two?

OpenStudy (salina):

can some one check my answer is 28

OpenStudy (anonymous):

you are computing \[\dbinom{10}{2}\] yes

OpenStudy (anonymous):

you get \[\frac{10\times 9}{2}=5\times 9=45\]

OpenStudy (anonymous):

not sure where the 28 came from oh yes i am. 28 is eight choose two. you want ten choose two

OpenStudy (anonymous):

not sure what notation you are using, but you want 10C2 which is the same as 10C8. either way. because of course if she makes 8 she has to miss 2 and vice versa.